# On Computational Limits and Provably Efficient Criteria of Visual Autoregressive Models: A Fine-Grained Complexity Analysis

### v1

@@
Yekun Ke keyekun0628@gmail.com. Independent Researcher.    Xiaoyu Li 7.xiaoyu.li@gmail.con. Independent Researcher.    Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhizhou Sha shazz20@mails.tsinghua.edu.cn. Tsinghua University.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at UC Berkeley.
Recently, Visual Autoregressive ($\mathsf{VAR}$) Models introduced a groundbreaking advancement in the field of image generation, offering a scalable approach through a coarse-to-fine “next-scale prediction” paradigm. However, the state-of-the-art algorithm of $\mathsf{VAR}$ models in [Tian, Jiang, Yuan, Peng and Wang, NeurIPS 2024] takes $O(n^{4})$ time, which is computationally inefficient. In this work, we analyze the computational limits and efficiency criteria of $\mathsf{VAR}$ Models through a fine-grained complexity lens. Our key contribution is identifying the conditions under which $\mathsf{VAR}$ computations can achieve sub-quadratic time complexity.
Specifically, we establish a critical threshold for the norm of input matrices used in $\mathsf{VAR}$ attention mechanisms. Above this threshold, assuming the Strong Exponential Time Hypothesis ($\mathsf{SETH}$) from fine-grained complexity theory, a sub-quartic time algorithm for $\mathsf{VAR}$ models is impossible. To substantiate our theoretical findings, we present efficient constructions leveraging low-rank approximations that align with the derived criteria.
@@
  • On the positive side, we show that when $d=O(\log n)$ and $R=o(\sqrt{\log n})$, there is an $O(n^{2+o(1)})$-time algorithm that approximates the output of $\mathsf{VAR}$ model up to $1/\operatorname{poly}(n)$ additive error.
  • On the negative side, we show that when $d=O(\log n)$ and $R=\Theta(\sqrt{\log n})$, assuming $\mathsf{SETH}$, it is impossible to approximate the output of $\mathsf{VAR}$ model up to $1/\operatorname{poly}(n)$ additive error in truly sub-quartic time $O(n^{4-\Omega(1)})$.
@@ 1 Introduction
Can we perform the computations of $\operatorname{\mathsf{VAR}}$ models faster than $O(n^{4})$ time?
@@ 1.1 Our Results
Our first result shows that when $R\geq\Omega(\sqrt{\log n})$, it is impossible to design a truly sub-quartic time algorithm. Our results for the lower bound make use of the Strong Exponential Time Hypothesis ($\mathsf{SETH}$[26] from the area of fine-grained complexity regarding the time required to solve $k$-SAT.
@@ 2.2 Acceleration via Low-rank Approximation
The low-rank approximation is a widely used technique for optimizing transformer training by reducing computational complexity [33, 37, 7, 24, 13, 34]. Specifically, [7] builds upon the low-rank approximation framework introduced in [5], which originally focused on forward attention computation, to approximate the gradient of attention mechanisms. This method effectively reduces the computational cost associated with gradient calculations. In [33], this low-rank gradient approximation approach is further extended to multi-layer transformers, demonstrating that backward computations in such architectures can be approximated in nearly linear time. Additionally, [37] generalizes the work of [7] to a tensor-based attention model, leveraging the forward computation results from [8] to enable efficient training of tensorized attention mechanisms. Finally, [24] utilizes low-rank approximation methods in the training process of Diffusion Transformers (DiTs)., highlighting the versatility of these methods in various transformer-based models.
@@ 3 Model Formulation
In Section 3.1, we give some definitions which will be used later. In Section 3.2, we present the mathematical formulation for the token map generation phase. In Section 3.3, we provide the mathematical formulation for the feature map reconstruction phase. Finally, in Section 3.4, we detail the mathematical formulation of the VQ-VAE Decoder within the $\operatorname{\mathsf{VAR}}$ model.
3.1 Notations and Definitions
We give the following notations in our setting. Given an integer $n\in\mathbb{Z}^{+}\cup\{0\}$, the set $\{1,2,\dots,n\}$ is represented by $[n]$. In our paper, nearly linear time is defined as $O(n\operatorname{poly}\log n)$, and almost linear time is defined as $O(n^{1+o(1)})$. Given a vector $c$, the diagonal matrix formed from $c$ is denoted as $\operatorname{diag}(c)$, where $c_{i}$ is the $i$-th diagonal entry of this matrix. Given a matrix $U$, we use $U^{\top}$ to denote the transpose of $U$. Given two vectors $a$ and $b$, which have the same length. The element-wise multiplication of $c$ and $d$ is denoted as $c\circ d$ with $i$-th entry being $c_{i}d_{i}$. Given a matrix $U$, we use $\|U\|_{F}$ to represent the Frobenius norm of $U$. Specifically, we have $\|U\|_{F}:=\sqrt{\sum_{i,j}U_{i,j}^{2}}$. Given a matrix $U$, we use $\|U\|_{\infty}$ to represent the maximum norm of $U$. Specifically, we have $\|U\|_{\infty}:=\max_{i,j}|U_{i,j}|$.
3.2 Phase 1: Token Maps Generation Phase
The $\operatorname{\mathsf{VAR}}$ model uses the $\operatorname{\mathsf{VAR}}$ Transformer to convert the initial tokens of the generated image into several pyramid-shaped token maps. And in the token maps generation phase, the token maps for the next scale, $M_{k+1}$, are generated based on the previous $k$ token maps ${M_{1},\dots,M_{k}}$. This phase has the main modules as the following:
@@ 3.2 Phase 1: Token Maps Generation Phase
Before inputting the $k$ token maps into the $\operatorname{\mathsf{VAR}}$ Transformer, the $\operatorname{\mathsf{VAR}}$ model upsamples the $(i)$-th token map to the size of the $(i+1)$-th token map. Specifically, the $\operatorname{\mathsf{VAR}}$ model employs an upsampling method using interpolation for the image. Here, we define the up-interpolation layers:
@@
Definition 3.2 (Up-interpolation Layer).
The Up-Interpolation layer is defined as follows:
  • Let $h\in\mathbb{N}$ and $h^{\prime}\in\mathbb{N}$ denote the height of the input feature map and output feature map, respectively.
  • Let $w\in\mathbb{N}$ and $w^{\prime}\in\mathbb{N}$ denote the width of the input feature map and the output feature map, respectively.
  • Let $c\in\mathbb{N}$ denote the number of channels of the input feature map and the output feature map.
  • Let $X\in\mathbb{R}^{h\times w\times c}$ denote the input feature map.
  • Let $Y\in\mathbb{R}^{h^{\prime}\times w^{\prime}\times c}$ denote the output feature map.
  • Let $s,t\in\{-1,0,1,2\}$.
    We use $\phi_{\rm up}:\mathbb{R}^{h\times w\times c}\to\mathbb{R}^{h^{\prime}\times w^{\prime}\times c}$ to denote the up-interpolation operation then we have $Y=\phi_{\rm up}(X)$. Specifically, for $i\in[h^{\prime}],j\in[w^{\prime}],l\in[c]$, we have
    $Y_{i,j,l}:=\sum_{s=-1}^{2}\sum_{t=-1}^{2}W(s)\cdot X_{\frac{ih}{h^{\prime}}+s,\frac{jw}{w^{\prime}}+t,l}\cdot W(t)$
Transformer Blocks.
After the up-sample process, the generated token maps above will be input into the Transformer to predict the next token map. Here, we define several blocks for the $\operatorname{\mathsf{VAR}}$ Transformer.
We first define the attention matrix.
Definition 3.3 (Attention Matrix).
Define the model weights as $W_{Q},W_{K}\in\mathbb{R}^{d\times d}$, and let $X\in\mathbb{R}^{n\times d}$ represent the input of length $n$. The attention matrix $A\in\mathbb{R}^{n\times n}$ is given by the following, for $i,j\in[n],$
$A_{i,j}:=$ $~{}\exp(X_{i,*}W_{Q}W_{K}^{\top}X_{j,*}^{\top}.$
To move on, we define the single attention layer in the following way:
@@
Let $X\in\mathbb{R}^{n\times d}$ be the input matrix. Let $W_{V}\in\mathbb{R}^{d\times d}$ represent the weighted matrix of value. Similar to the standard attention mechanism, the goal is to produce an $n\times d$ output matrix, where $D:=\operatorname{diag}(A{\bf 1}_{n})\in\mathbb{R}^{n\times n}$. We define attention layer $\mathsf{Attn}$ as the following:
$\mathsf{Attn}(X):=$ $~{}D^{-1}AXW_{V}.$
Firstly, we give the definition of the $\operatorname{\mathsf{VAR}}$ Transformer Layer.
Definition 3.5 ($\operatorname{\mathsf{VAR}}$ Transformer Layer).
Given that the following conditions are true:
  • Define $X\in\mathbb{R}^{n\times d}$ as the input data matrix.
  • Let the up-interpolation layer $\phi_{\mathrm{up}}$ as given in Definition 3.2.
  • Let the $\mathsf{Attn}(X)$ as given in Definition 3.3.
Then, we defined one $\operatorname{\mathsf{VAR}}$ Transformer Layer as
$\mathsf{F}_{\mathrm{var}}(X):=\mathsf{Attn}(\phi_{\rm up}(X))$
Then, we can give the definition of the Fast $\operatorname{\mathsf{VAR}}$ Transformer Layer.
Definition 3.6 (Fast $\operatorname{\mathsf{VAR}}$ Transformer Layer).
Given that the following conditions are true:
  • Let $X\in\mathbb{R}^{n\times d}$ denote the input data matrix.
  • Let the up-interpolation layer $\phi_{\rm up}$ as given in Definition 3.2.
  • Let the $\mathsf{AAttC}(X)$ as given in Definition 4.2.
Then, we defined one Fast $\operatorname{\mathsf{VAR}}$ Transformer Layer as
$\mathsf{F}_{\mathrm{fvar}}(X):=\mathsf{AAttC}(\phi_{\rm up}(X))$
Definition 3.7 ($\operatorname{\mathsf{VAR}}$ Model).
Given that the following conditions are true:
  • Let the initial input be defined as $x_{1}\in\mathbb{R}^{1\times d}$.
  • Let the $\operatorname{\mathsf{VAR}}$ Transformer layer $\mathsf{F}_{\mathrm{var}}$ as given in Definition 3.5.
  • Let $r\in\mathcal{N}$ represent the total iteration of $\operatorname{\mathsf{VAR}}$ model.
Then, for $i\in\{0,1,2,\cdots,r\}$, we defined the intermediate variable $T_{i}(X)$ as follows:
$T_{i}(X)=\begin{cases}x_{1},&~{}i=0\\ \mathsf{F}_{\mathrm{var}}(T_{i-1}(X)),&~{}i\in[r]\\ \end{cases}$
Then, the final output of the $\operatorname{\mathsf{VAR}}$ model is $T_{r}(X)$.
Definition 3.8 (Fast $\operatorname{\mathsf{VAR}}$ Model).
Given that the following conditions are true:
  • Define $x_{1}\in\mathbb{R}^{1\times d}$ as the initial input.
  • Let the Fast $\operatorname{\mathsf{VAR}}$ Transformer layer $\mathsf{F}_{\mathrm{fvar}}$ be defined as Definition 3.6.
  • We take $R$ to be the upper bound on the entries of the input matrices involved in attention computations for $\operatorname{\mathsf{VAR}}$ models.
  • Let $r\in\mathcal{N}$ denote the total iteration of the Fast $\operatorname{\mathsf{VAR}}$ model.
Then, for $i\in\{0,1,2,\cdots,r\}$, we defined the intermediate variable $\tilde{T}_{i}(X)$ as follows:
$\widetilde{T}_{i}(X)=\begin{cases}x_{1},&~{}i=0\\ \mathsf{F}_{\mathrm{fvar}}(\widetilde{T}_{i-1}(X)),&~{}i\in[r]\\ \end{cases}$
Then, the final output of the $\operatorname{\mathsf{VAR}}$ model is $\tilde{T}_{r}(X)$.
3.3 Phase 2: Feature Map Reconstruction
In phase 2, the $\operatorname{\mathsf{VAR}}$ model will transform the generated token maps into feature maps. This phase has the following main modules:
@@ 3.3 Phase 2: Feature Map Reconstruction
  • Let $X\in\mathbb{R}^{h\times w\times c_{\rm in}}$ represent the input feature map.
  • For $l\in[c_{\rm out}]$, we use $K^{l}\in\mathbb{R}^{3\times 3\times c_{\rm in}}$ to denote the $l$-th convolution kernel.
@@
  • Let $Y\in\mathbb{R}^{h\times w\times c_{\rm out}}$ represent the output feature map.
We use $\phi_{\rm conv}:\mathbb{R}^{h\times w\times c_{\rm in}}\to\mathbb{R}^{h\times w\times c_{\rm out}}$ to represent the convolution operation then we have $Y=\phi_{\rm conv}(X)$. Specifically, for $i\in[h],j\in[w],l\in[c_{\rm out}]$, we have
$Y_{i,j,l}:=\sum_{m=1}^{3}\sum_{n=1}^{3}\sum_{c=1}^{c_{\rm in}}X_{i+m-1,j+n-1,c}\cdot K^{l}_{m,n,c}+b$
@@
3.4 Phase 3: VQ-VAE Decoder process

### v2

@@
Yekun Ke keyekun0628@gmail.com. Independent Researcher.    Xiaoyu Li xiaoyu.li2@student.unsw.edu.au. University of New South Wales.    Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhizhou Sha shazz20@mails.tsinghua.edu.cn. Tsinghua University.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at UC Berkeley.
Recently, Visual Autoregressive ($\mathsf{VAR}$) Models introduced a groundbreaking advancement in the field of image generation, offering a scalable approach through a coarse-to-fine “next-scale prediction” paradigm. Suppose that $n$ represents the height and width of the last VQ code map generated by $\mathsf{VAR}$ models, the state-of-the-art algorithm in [Tian, Jiang, Yuan, Peng and Wang, NeurIPS 2024] takes $O(n^{4+o(1)})$ time, which is computationally inefficient. In this work, we analyze the computational limits and efficiency criteria of $\mathsf{VAR}$ Models through a fine-grained complexity lens. Our key contribution is identifying the conditions under which $\mathsf{VAR}$ computations can achieve sub-quadratic time complexity. We have proved that assuming the Strong Exponential Time Hypothesis ($\mathsf{SETH}$) from fine-grained complexity theory, a sub-quartic time algorithm for $\mathsf{VAR}$ models is impossible. To substantiate our theoretical findings, we present efficient constructions leveraging low-rank approximations that align with the derived criteria.
@@
  • On the negative side, we show that when $d=O(\log n)$ and $R=\Theta(\sqrt{\log n})$, assuming $\mathsf{SETH}$, it is impossible to approximate the output of $\mathsf{VAR}$ model up to $1/\mathrm{poly}(n)$ additive error in truly sub-quartic time $O(n^{4-\Omega(1)})$.
  • On the positive side, we demonstrate a sharp phase transition in the computational complexity of the $\mathsf{VAR}$ model. When $d=O(\log n)$ and $R=o(\sqrt{\log n})$, the runtime improves dramatically from $O(n^{4+o(1)})$ to $O(n^{2+o(1)})$, which approximates the output of the output of $\mathsf{VAR}$ model up to $1/\mathrm{poly}(n)$ additive error.
@@ 1 Introduction
Can we perform the computations of $\operatorname{\mathsf{VAR}}$ models
faster than $O(n^{4})$ time?
@@ 1.1 Our Results
Our first result shows that when the attention entry range $R\geq\Omega(\sqrt{\log n})$, it is impossible to design a truly sub-quartic time algorithm. Our results for the lower bound make use of the Strong Exponential Time Hypothesis ($\mathsf{SETH}$[26] from the area of fine-grained complexity regarding the time required to solve $k$-SAT.
@@ 2.2 Acceleration via Low-rank Approximation
The low-rank approximation is a widely used technique for optimizing transformer training by reducing computational complexity [34, 38, 7, 24, 13, 35, 30]. Specifically, [7] builds upon the low-rank approximation framework introduced in [5], which originally focused on forward attention computation, to approximate the gradient of attention mechanisms. This method effectively reduces the computational cost associated with gradient calculations. In [34], this low-rank gradient approximation approach is further extended to multi-layer transformers, demonstrating that backward computations in such architectures can be approximated in nearly linear time. Additionally, [38] generalizes the work of [7] to a tensor-based attention model, leveraging the forward computation results from [8] to enable efficient training of tensorized attention mechanisms. Finally, [24] utilizes low-rank approximation methods in the training process of Diffusion Transformers (DiTs)., highlighting the versatility of these methods in various transformer-based models.
@@ 3 Model Formulation
In this section, we begin by introducing the notations used throughout the paper in Section 3.1. Section 3.2 presents the overall architecture of the VAR model and divides its processing workflow into three stages. In Section 3.3, we provide the mathematical formulation for the modules involved in the pyramid-shaped token map generation stage. Section 3.4 offers the mathematical formulation for the modules in the feature map reconstruction stage, while Section 3.5 presents the mathematical formulation for the modules in the VQ-VAE Decoder process stage.
3.1 Notations
Given an integer $n\in\mathbb{Z}^{+}$, we use $[n]$ to denote the set $\{1,\dots,n\}$. Given a vector $c$, the diagonal matrix formed by $c$ is denoted as $\operatorname{diag}(c)$, where $c_{i}$ denotes the $i$-th diagonal element. Given a matrix $U$, we use $U^{\top}$ to denote the transpose of $U$. Given a matrix $U$, we denote it Frobenius norm as $\|U\|_{F}$, which is defined as $\|U\|_{F}:=\sqrt{\sum_{i,j}U^{2}_{i,j}}$. Additionally, we define $\|U\|_{\infty}$ as the maximum norm of $U$, which is defined as $\|U\|_{\infty}=\max_{i,j}|U_{i,j}|$. Given two vectors $c,d\in\mathbb{R}^{n}$, the notation $c\circ d$ represents the element-wise product (also known as the Hadamard Product). It is defined as: $c\circ d=(c_{1}d_{1},\dots,c_{n}d_{n})$. In our paper, nearly linear time is defined as $O(n\operatorname{poly}\log n)$, and almost linear time is defined as $O(n^{1+o(1)})$.
3.2 Overall Architecture
In this section, We present the overall architecture of the $\operatorname{\mathsf{VAR}}$ model and divide its processing workflow into three stages.
Stage 1: Pyramid-Shaped Token Maps Generation. Firstly, the $\operatorname{\mathsf{VAR}}$ model will start by quantizing an initial input token map $X_{\mathrm{init}}\in\mathbb{R}^{1\times 1\times d}$ into $K$ multiple scale pyramid-shaped token maps $(r_{1},\dots,r_{K})$, each at an increasingly higher resolution $h_{k}\times w_{k}$. During the $k$-th autoregressive step, all the $h_{k}\times w_{k}$ will be generated in parallel, conditioned on $r_{k}$’s prefix $r_{1},\dots,r_{k-1}$. In Section 3.3, we provide a mathematical definition for each module in this stage.
Stage 2: Feature Map Reconstruction. The second stage of the $\operatorname{\mathsf{VAR}}$ model is to reconstruct the generated pyramid-shaped token maps $r_{1},\dots,r_{K}$ into a Feature Map. Specifically, the $\operatorname{\mathsf{VAR}}$ model uses an up-interpolation layer to interpolate each of the token maps $(r_{1},...,r_{K-1})$ to the size of $r_{K}$ and applies a convolution layer to reduce the loss introduced by the interpolation. After this process, the VAR model sums the K token maps to obtain the desired Feature Map. In Section 3.4, we provide a mathematical definition for each module in this stage.
Stage 3: Generating Image Using VQ-VAE Decoder. The third stage of $\operatorname{\mathsf{VAR}}$ model is to use VQ-VAE Decoder to generate the final output image by taking the input of feature map. We follow the implementation of [52] and regard the VQ-VAE Decoder as a module composed of fixed-depth ResNet layers, attention layers, and up-interpolation layers. In Section 3.5, we provide a mathematical definition for each module in this stage.
3.3 Stage 1: Token Maps Generation
The $\operatorname{\mathsf{VAR}}$ model uses the $\operatorname{\mathsf{VAR}}$ Transformer to convert the initialized token map $X_{\mathrm{init}}$ into a series of pyramid-shaped token maps. The $\operatorname{\mathsf{VAR}}$ Transformer alternates between up sample blocks and attention layers to get the output.
@@ 3.3 Stage 1: Token Maps Generation
The $k$-th up sample block takes as input the initial token map $X_{\mathrm{init}}$ and the previous pyramid-shaped token maps $X_{1},\ldots,X_{k}$, sets $Y_{1}=X_{\mathrm{init}}$ and up samples each $X_{i}$ into a new token map $Y_{i+1}$, and outputs the new pyramid-shaped token maps $Y_{1},\ldots,Y_{k+1}$.
The upsampling on each token map $X_{r}(r\in[k])$ uses interpolation with a bicubic spline kernel.
@@
Definition 3.2 (Up-interpolation Layer for One-Step Geometric Sequence).
The layer $\phi_{{\rm up},r}$ takes the input feature map $X_{r}\in\mathbb{R}^{h_{r}\times w_{r}\times d}$ and computes the output feature map $Y_{r+1}\in\mathbb{R}^{h_{r+1}\times w_{r+1}\times d}$, where $h_{r}<h_{r+1}$ are the heights, $w_{r}<w_{r+1}$ are the widths, and $d\in\mathbb{N}$ is the hidden dimension. It computes $Y_{r+1}=\phi_{{\rm up},r}(X_{r})$ with a bicubic spline kernel $W$: for $i\in[h_{r+1}],j\in[w_{r+1}],l\in[d]$,
$[Y_{r+1}]_{i,j,l}:=\sum_{s=-1}^{2}\sum_{t=-1}^{2}W(s)\cdot[X_{r}]_{\frac{i\cdot h_{r}}{h_{r+1}}+s,\frac{j\cdot w_{r}}{w_{r+1}}+t,l}\cdot W(t)$ (1)
Figure 1: Example of the Pyramid Up-Interpolation Layer $\Phi_{{\rm up},2}$ used in the model.
We are now ready to present the up sample block $\Phi$.
Definition 3.3 (Pyramid Up-Interpolation Layer $\Phi$).
The layer $\Phi_{{\rm up},k}$ takes the initial token map $X_{\mathrm{init}}$ and the token maps $X_{r}\in\mathbb{R}^{h_{r}\times w_{r}\times c}(r\in[k])$ and computes new token maps $Y_{r}\in\mathbb{R}^{h_{r}\times w_{r}\times c}$. It sets $Y_{1}=X_{\mathrm{init}}$ and computes $Y_{r+1}=\phi_{{\rm up},r}(X_{r})$ as in Definition 3.2. The output is the set consisting $Y_{i}(i\in[k+1])$.
Attention Layer.
After an up sample block, the token maps (after being flattened into a proper shape) will be input into an attention layer.
@@
Let $X\in\mathbb{R}^{n\times d}$ denote the input matrix. Let $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d\times d}$ denote the weight matrix for query, key, and value, respectively. First, compute the attention matrix $A\in\mathbb{R}^{n\times n}$:
[…]
Then, compute the output:
$\mathsf{Attn}(X):=$ $~{}D^{-1}AXW_{V},$
where $D:=\operatorname{diag}(A{\bf 1}_{n})\in\mathbb{R}^{n\times n}$
$\operatorname{\mathsf{VAR}}$ Transformer.
A $\operatorname{\mathsf{VAR}}$ Transformer with $K$ layers alternates between the attention layer and up sample blocks (where the output of each layer is reshaped to a proper shape as the input for the next layer):
Definition 3.5 ($\operatorname{\mathsf{VAR}}$ transformer).
The transformer $\mathsf{TF}$ takes an initial token map $X_{\mathrm{init}}\in\mathbb{R}^{1\times d}$, computes
  • $Z_{0}=X_{\mathrm{init}}$,
[…]
and finally outputs $\mathsf{Attn}_{K}(Z_{K-1})$. Here $\Phi_{\mathrm{up},k}$ is defined in Definition 3.3, $\mathsf{Attn}_{i}$ is defined in Definition 3.4, $Z_{k-1}$ is flatten into shape $(\sum_{r=1}^{k}h_{r}w_{r})\times d$ as input for $\mathsf{Attn}_{k}$, and the output of $\mathsf{Attn}_{k}$ is reshaped into $X_{r}\in\mathbb{R}^{h_{r}\times w_{r}\times c}(r\in[k])$ as input for $\Phi_{\mathrm{up},k}$.
For convenience, we often abuse notation slightly and write:
$\mathsf{TF}(X_{\mathrm{init}}):=$ $~{}\mathsf{Attn}_{K}\circ\Phi_{\mathrm{up},K-1}\circ\cdots\circ\Phi_{\mathrm{up},1}\circ\mathsf{Attn}_{1}(X_{\mathrm{init}}),$
where $\circ$ denotes function composition.
3.4 Stage 2: Feature Map Reconstruction
In phase 2, the $\operatorname{\mathsf{VAR}}$ model will transform the generated pyramid-shaped token maps into feature maps. This phase has the following main modules:
@@ 3.4 Stage 2: Feature Map Reconstruction
  • Let $X\in\mathbb{R}_{p}^{h\times w\times c_{\rm in}}$ represent the input feature map.
  • For $l\in[c_{\rm out}]$, we use $K^{l}\in\mathsf{F}_{p}^{3\times 3\times c_{\rm in}}$ to denote the $l$-th convolution kernel.
@@
  • Let $Y\in\mathbb{R}_{p}^{h\times w\times c_{\rm out}}$ represent the output feature map.
We use $\phi_{\rm conv}:\mathbb{R}_{p}^{h\times w\times c_{\rm in}}\to\mathbb{R}_{p}^{h\times w\times c_{\rm out}}$ to represent the convolution operation then we have $Y=\phi_{\rm conv}(X)$. Specifically, for $i\in[h],j\in[w],l\in[c_{\rm out}]$, we have
$Y_{i,j,l}:=\sum_{m=1}^{3}\sum_{n=1}^{3}\sum_{c=1}^{c_{\rm in}}X_{i+m-2,j+n-2,c}\cdot K^{l}_{m,n,c}+b$
@@
3.5 Stage 3: VQ-VAE Decoder Process